A poker iron has one end in a fire of temperature T = 500° C. The other end is at 26° C. Its length is 1.2 m and is radius is 5 x 10-3 m. How much heat, Q, is conducted from one end of the poker to the other in 10 s? kiron = 79 (J/s∙m∙°C)
1
Expert's answer
2020-04-28T09:28:45-0400

The Fourier law says

"Q=\kappa\frac{\Delta T}{\Delta h}A\tau"

"=79\times \frac{500-26}{1.2}\times 3.14\times (5\times 10^{-3})^2\times 10"

"=24.5\:\rm J."